A Marvel comic character fires a missile, towards a DC Comic hero, the missile gets splitted into three fragments. Two fragments each of mass 100 kg, move away from each other making an angle of 120", with a speed of 200 m/s. Find the velocity of the last fragment having, mass of 200 kg. Also calculate Energy released in the process.

1
Expert's answer
2022-03-06T18:05:50-0500

"Mu=2mv\cos\frac \alpha 2,"

"u=\frac{2mv\cos\frac \alpha 2}M=100~\frac ms,"

"E=mv^2+\frac{Mu^2}2=3~MJ."